Probability that A speaks truth is $\frac{4}{5}$ . A coin is tossed. A reports that a head appears. The probability that actually there was head is:
$\frac{4}{5}$
Let H denote the event that the coin shows a head and T that it shows a tail, with \(P(H)=P(T)=1/2\) for a fair coin.
\(P(\text{reports head}\mid H) = 4/5\) (A speaks the truth), and \(P(\text{reports head}\mid T) = 1-4/5 = 1/5\) (A lies).
By Bayes' theorem, \(P(H\mid\text{reports head}) = \dfrac{P(H)P(\text{reports head}\mid H)}{P(H)P(\text{reports head}\mid H)+P(T)P(\text{reports head}\mid T)}\).
\(= \dfrac{(1/2)(4/5)}{(1/2)(4/5)+(1/2)(1/5)} = \dfrac{2/5}{2/5+1/10} = \dfrac{2/5}{1/2} = \dfrac{4}{5}\).
Bag I contains 3 Red and 4 Black Balls and Bag II contains 4 Red and 5 Black Balls. One ball is transferred from Bag I to Bag II and then a ball drawn from Bag II. The ball so drawn is found to be red in colour. Find the probability that the transferred ball is black.
One bag contain 3 white and 2 black balls, another bag contains 5 white and 3 black balls. If a bag is chosen at random and a ball is drawn from it, what is the chance that it is white?
Three cooks X, Y and Z bake a special kind of cake, and with respective probabilities 0.02, 0.03 and 0.05, it fails to rise. In the restaurant where they work, X bakes 50%, Y bakes 30% and Z bakes 20% of cakes. What is the proportion of failures caused by X?
An insurance company insured 2000 scooter drivers, 4000 car drivers and 6000 truck drivers. The probabilities of an accident involving a scooter driver car driver and a truck driver are 0.01, 0.03 and 0.15 respectively. One of the insured persons meets with an accident. The probability that the person is a scooter driver is
There are 3 coins in a box. One is a two-headed coin; another is a fair coin; and third is biased coin that comes up heads 75% of time. When one of the three coins is selected at random and flipped, it shows heads. What is the probability that it was the two-headed coin?
In a bolt factory, machines X, Y, Z manufacture bolts that are respectively 25% 35%, and 40% of the factory’s total output. The machines X, Y, Z respectively produce 2%, 4% and 5% defective bolts. A bolt is drawn at random from the product and is found to be defective. What is the probability that it was manufactured by machine X?